OpenStudy (anonymous):

Write an equation in the form y = mx + b for the line that passes through the point (-5, -1) and has a slope of -2.

jimthompson5910 (jim_thompson5910):

y = mx + b y = -2x + b -1 = -2(-5) + b -1 = 10 + b Keep going to solve for b